# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ III-C2 Observation Viewpoint of the World
Robots acquire skills by observing and imitating humans or other robots in their environment. Depending on the observation viewpoint, world models for robot learning can be categorized into third-person (exocentric) [21, 25, 26] and first-person (egocentric) [27, 28] perspectives. Many existing methods learn from exocentric perspectives, capturing skills from an external viewpoint [21, 25, 26]. However, exocentric observations do not fully align with how humans perceive the world. This has motivated the development of egocentric world models. For example, Chen et al. [27] observe a continuous loop of human interactions, in which humans perceive egocentric observations and take 3D actions repeatedly. They model these interactions as sequences of “state-action-state-action” tokens, processed using a causal attention mechanism. Zhang et al. [7] focus on multi-agent planning, inferring other agents’ actions from world states estimated via partial egocentric observations.
Grauman et al. [28] argue that egocentric and exocentric viewpoints are complementary. Learning through egocentric viewpoints allows robots to better understand hand-object interactions and the attention mechanism of the camera wearer, while exocentric perspectives provide information about the surrounding environment and whole-body poses.
III-C3 Representation of the World.
A central aspect of world models lies in how the environment is represented, which directly influences their ability to reason about dynamics, predict future states, and generalize across tasks. World representations can be broadly categorized into scene-centric, object-centric, and flow-centric approaches. In scene-centric representations, the environment is encoded as a single holistic latent, typically learned directly from pixels or raw sensory inputs [29, 12, 13, 15, 100]. While video generation tasks aim to maximize the visual fidelity of predicted sequences, robotic manipulation often does not require the full visual detail. Irrelevant elements such as the background or parts of the robot arm can be ignored. This motivates the use of object-centric representations, which focus on task-relevant entities and their interactions [25, 33, 26, 69, 18]. Flow-centric representations, in contrast, are designed to capture the motion dynamics of the environment, emphasizing temporal change and spatial displacement [65].
@@ III-D Task Scope
World models can also be categorized based on their task coverage. Some studies focus on single-task objectives, such as future-scene prediction [37, 36, 33, 38, 30], or planning and action prediction [39].
In contrast, an increasing number of studies aim to support multiple tasks simultaneously, thereby enhancing the generality and applicability of world models. For instance, Cheang et al. [58, 66, 5, 65] generate videos for future-scene prediction and accordingly infer corresponding actions. Other works pursue simultaneous action prediction and world-scene forecasting [40, 27, 41, 42]. Beyond dual-task integration, several approaches extend world models to even broader capabilities. For instance, Bruce et al. [20] propose interactive video generation that supports environment prediction and imitation learning, and utilize a latent action model to infer policies from unseen, action-free videos. Liao et al. [57] introduce a unified framework for embodied video generation, policy learning, and simulation. Lu et al. [34] learn 3D world representations for future-state prediction, imitation learning, and simulator through video generation. Zhu et al. [43] develop an action-conditioned world model supporting trajectory-conditioned video generation, policy evaluation, and planning. Similarly, Huang et al. [8] achieve multi-view video generation, robotic action prediction, and a data flywheel mechanism for sim-to-real adaptation.
Would foundation models. When discussing task scope, the notion of “foundation world models” becomes essential. These approaches aim to generalize across diverse tasks through large-scale training, paving the way for world models that act as universal backbones for robotics. One line of research achieves this through large-scale pretraining followed by task-specific fine-tuning [34, 6, 96, 58, 101]. In particular, Mazzaglia et al. [101] integrate a foundation VLM with a generative world model to enhance multimodal generalization. Other works directly pursue large-scale end-to-end training to build general-purpose world models [20, 40].
@@
IV Functions of the World Model
World models play a central role in modern robotics by providing an internal predictive understanding of the environment. They enable robots to reason about future states, anticipate the consequences of actions, and perform counterfactual evaluations, which are particularly valuable in real-world settings where interactions are costly, risky, or time-consuming. By modeling environmental dynamics, world models form the foundation for autonomous, adaptable, and efficient robotic systems. In robotics, world models serve two complementary functions: decision support, by predicting future scenes, actions and planning, and training facilitation, by generating data or acting as learned simulators. These roles are often closely related. For example, a world model used as a simulator can simultaneously generate training data and assist decision making [34, 57]. By combining these functionalities, world models provide a comprehensive framework that enables robots to act intelligently, learn efficiently, and adapt to complex and dynamic environments. Additional details of the world models are provided in Table I, which complements the following discussion.
@@ IV-A1 Implicit World Models for Action Prediction and Planning
This line of work explores world models that enable action prediction and planning without explicitly modeling state transitions or world dynamics. These approaches typically leverage the strong reasoning and next-token prediction capabilities of Large Language Models (LLMs), Vision-Language Models (VLMs), and Vision-Language-Action (VLA) models. Since LLMs lack direct access to environmental or robotic states, auxiliary components are often incorporated to provide grounding. For example, Ahn et al. [62] introduce affordance functions to evaluate the feasibility of skills for completing a target task. Xiang et al. [72, 54] employ encoders to process environmental information, while Zhang et al. [60] integrate multimodal tokens including states, images, and text to enhance reasoning and generalization. Zhang et al. [17] further combine 2D and 3D encoders to process RGB images and 3D point clouds, capturing complementary spatial cues for richer world understanding. Hong et al. [18] extend this paradigm by incorporating additional sensory modalities such as vision, audio, tactile, and thermal inputs to achieve a more comprehensive understanding of the environment.
@@
Conventional LLMs are language-centric and typically treat visual and other sensory information as auxiliary inputs. VLMs extend this paradigm by jointly learning aligned visual and linguistic representations, enabling grounded perceptual understanding of the world [83]. Zhang et al. [84] further leverage VLMs to generate candidate action sequences, which are evaluated using a lightweight action-conditioned video prediction model to forecast future scenes. The predicted outcomes are then assessed by the VLM to select the final action. An increasing number of studies extend VLMs to VLA by equipping them with low-level action generation capabilities. For instance, Zitkovich et al. [16] represent robot actions as a form of language, effectively bridging perception and control through textual grounding. Zhen et al. [41] employ a 3D-based LLM [85] to represent and predict 3D world states and generate actions, incorporating a diffusion model to synthesize future scenes. Inspired by the dual-process theory of human cognition [102], Björck et al. [44] design a dual-system architecture in which a VLM serves as the reasoning module (System 2) and a Diffusion Transformer functions as the action module (System 1), with both components jointly optimized for coordinated reasoning and actuation. Zhou et al. [103] preserve the reasoning capability of VLMs while introducing a Mixture-of-Experts (MoE) to alleviate conflicts between multimodal understanding and robotic manipulation in the parameter space. Kim et al. [55] train their model on a large corpus of real-world robot demonstrations, enabling efficient adaptation to new robotic platforms through parameter-efficient fine-tuning.
@@
To further enhance long-horizon prediction, reasoning, and imagination, several methods integrate large language or multimodal models into other world model architectures, where they serve as core components. For instance, Chen et al. [27] employ the open-source LLM, i.e., InternLM [87], to predict future states from egocentric observations as a fundamental element of the JEPA framework. Similarly, Vision-Language Models [101] and Video-Language Models [86] have been incorporated into Dreamer-style architectures for low-level dynamics modeling, where they extract high-level semantic knowledge of the world to guide prediction.
@@
Notably, LLMs, VLMs, and VLAs can also act as explicit world models that predict future scenes [78, 60, 79] or world knowledge [60]. We will elaborate them in Section IV-A3.
@@ IV-A2 Latent Dynamics Modeling for Action Prediction and Planning
This line of research focuses on modeling the temporal evolution of environment dynamics within a latent space, facilitating efficient action prediction, planning and future imagination. Operating in a compact latent space requires fewer environment interactions and reduces computational cost compared to pixel-based modeling. Hafner et al. [29, 12, 13, 14, 15] introduce online planning in latent space through the Recurrent State-Space Model (RSSM), which learns to reconstruct input observations [29]. The Dreamer series [12, 13, 14, 15] introduces latent imagination, allowing agents to predict and plan over latent trajectories instead of pixels for more efficient and stable policy learning. Specifically, DreamerV1 [12] learns long-horizon behaviors from images by jointly predicting actions and state values, greatly improving sample efficiency. DreamerV2 [13] extends this framework to discrete environments by introducing binary latent variables, achieving human-level performance on the Atari benchmark. DreamerV3 [15, 100] further improves scalability and generative capacity through techniques such as symlog normalization for reward stabilization, refined KL balancing, and enhanced replay buffers. Sekar et al. [39] enhance generalization to downstream tasks through self-supervised learning without task-specific rewards, while Wu et al. [14] deploy Dreamer in the real world without simulators. Gumbsch et al. [56] introduce context-sensitive dynamics via a context-specific RSSM and hierarchical architecture to improve scalability and long-horizon prediction. Ferraro et al. [25] develop object-centric world models for improved interaction reasoning.
Under the JEPA framework, Chen et al. [27] capture causal and temporal dependencies by organizing states and actions into an interleaved sequence, integrating future state prediction and action generation within a unified transformer architecture. Building on this, Assran et al. [77] leverage pre-trained video encoders optimized with a masked denoising objective as the core of JEPA, enabling self-supervised learning through an action-conditioned predictor that autoregressively forecasts future states and actions. Incorporating other potential world models, such as LLMs and VLMs, have been introduced in Section IV-A1.
There are also approaches that couple Model Predictive Control (MPC) with learned world models, where the predictive model is used to simulate future trajectories and select optimal actions in a receding-horizon manner. For example, Hansen et al. [104] learn task-specific latent dynamics models using temporal-difference objectives and apply them for efficient online Model Predictive Control. Hansen et al. [105] further improve generalization across diverse embodiments and action spaces by learning an implicit, control-centric dynamics model.
@@ IV-A3 Vision-based Action Prediction and Planning
Vision-Conditioned Goals. Finn et al. [36] learn to predict motion dynamics that remain consistent across visual appearances, aiming to enable long-range, action-conditioned video prediction and generalization to unseen objects. Ebert et al. [30, 64] improve long-horizon prediction by using an image registration–based cost function that continuously corrects errors during execution, achieving closed-loop visual planning. Bu et al. [22] further extend this idea with text-conditioned video generation to synthesize depth- and flow-consistent sub-goal images. A feedback mechanism then selects sub-goals and generates corresponding actions based on visual error evaluation, bridging visual planning and policy learning.
@@
Imagining the future does not inherently produce actions. To enable action predictions. Finn et al. [63, 30, 64] incorporate visual prediction models with model-predictive control (MPC) to select the best action (sequence). Bu et al. [22] use an error-measurement strategy to select the best sub-goal images and an MLP that is train with an Inverse Dynamics objective to decode the corresponding actions.
@@
Language-Conditioned Goals. In [36, 63, 30, 64], task specifications are provided as goal images, which are often difficult to obtain and prone to over- or under-specification. To address this limitation, a growing line of research leverages language as a more flexible, compact, and general medium for specifying tasks. However, translating language instructions into precise, actionable representations grounded in the robot’s observations remains challenging due to the misalignment between linguistic descriptions and visual perception. To bridge this gap, Nair et al. [76] use action-conditioned video prediction to simulate future scenes under different action sequences and learn a language-conditioned reward function from crowd-sourced descriptions to measure task completion. The best sequence is selected to maximize the reward. Zhang et al. [78] take advantage of the semantic knowledge and reasoning abilities of VLA and incorporate a decoder into VLA to enable future scene predictions and action generation. Zhou et al. [66] parse language instructions into compositional primitives to capture spatial object relations and generalize to novel commands, while also supporting multimodal task inputs such as goal images and sketches. Zhang et al. [60] enhance reasoning and generalization by introducing additional dream queries that capture historical information and predict dynamic regions, depth, and semantic maps using foundation models such as DINOv2 [106] and SAM [107].
@@
Diverse Goals. Some works leverage diverse goal conditions to improve task understanding and completion. For instance, Wang et al. [9] develop a language–gesture-conditioned video generation model to disambiguate task specifications and integrate a behavior-cloning policy that unifies visual plan generation and manipulation. Du et al. [59] incorporate observed images as additional context in each frame-denoising step to synthesize video plans and employ an inverse-dynamics model (IDM) to infer the corresponding action sequences. Zhao et al. [79] introduce visual chain-of-thought (CoT) reasoning into VLA models by autoregressively generating sub-goal images alongside language instructions, enabling temporal planning and improving reasoning capability.
@@
Action inference. A key advantage of vision–based action prediction is that it does not rely on large-scale action-labeled data. They can be pre-trained on large-scale video data and infer actions by training a simple action extractor using small amounts of action data. Techniques for action extraction include the inverse dynamics model [108, 59, 66], a transformer encoder-decoder architecture [9]. Zhang et al. [7] use vision language models to propose actions, and a tree search to find the best plan. However, video predictions would contain irrelevant information to the target tasks or actions to execute such as background and robot arm. To handle this, Zhi et al. [69] extract 3D Flow from video data and learn 3D optical flow as a representation of object motions to guide action planning. Zhang et al. [60] propose dynamic region-based forecasting, which leverages optical flow prediction model [109, 110] to identify dynamic regions within the scene, enabling the model to concentrate on areas of motion that are critical for task execution instead of redundant frame reconstruction. Agarwal et al. [96] leverage large-scale pre-training on both image and post-training for robotic manipulation, including instruction-based video prediction and action-based next-frame prediction. 3D positional embeddings, including 3D factorized Rotary Position Embedding (RoPE) and absolute positional embedding (APE) for relative positions and absolute coordinates respectively, are adopted to capture spatial and temporal information. Actions are predicted through an action embedder MLP. Tian et al. [61] propose an end-to-end Predictive Inverse Dynamics Models (PIDM), which learn actions and visual futures synergistically to enhance the simulation and action predictions ability. [111] predict both future frames and robot actions within joint latent denoising process, which support planning and acting in a closed-loop manner.
@@
Visual Fidelity vs. Action Prediction. Guo et al. [21] hypothesize that models trained solely with frame-prediction losses tend to emphasize visual appearance fidelity while underestimating accurate dynamics modeling. This highlights the need for approaches that explicitly separate dynamics learning from visual rendering. To address this, FlowDreamer adopts a two-stage framework that first predicts environment dynamics and then renders corresponding visual observations.

### v2

@@ III-C2 Observation Viewpoint of the World
Robots acquire skills by observing and imitating humans or other agents in their environment. Depending on the observation viewpoint, world models for robot learning can be categorized into third-person (exocentric) [guo2025flowdreamer, ferraro2025focus, villar2025playslot] and first-person (egocentric) [chen2025egoagent, grauman2024ego] perspectives. Many existing methods learn from exocentric perspectives, capturing skills from an external viewpoint [guo2025flowdreamer, ferraro2025focus, villar2025playslot]. However, exocentric observations do not fully align with how humans perceive the world. This has motivated the development of egocentric world models. For example, Chen et al. [chen2025egoagent] observe a continuous loop of human interactions, in which humans perceive egocentric observations and take 3D actions repeatedly. They model interactions as sequences of “state-action-state-action” tokens, processed using a causal attention mechanism. Zhang et al. [zhang2025combo] focus on multi-agent planning, inferring other agents’ actions from world states estimated via partial egocentric observations.
Grauman et al. [grauman2024ego] argue that egocentric and exocentric viewpoints are complementary. Learning through egocentric viewpoints allows robots to better understand hand-object interactions and the attention mechanism of the camera wearer, while exocentric perspectives provide information about the surrounding environment and whole-body poses.
III-C3 Representation of the World
A central aspect of world models lies in how the environment is represented, which directly influences their ability to reason about dynamics, predict future states, and generalize across tasks. World representations can be broadly categorized into scene-centric, object-centric, and flow-centric approaches. In scene-centric representations, the environment is encoded as a single holistic latent, typically learned directly from pixels or raw sensory inputs [hafner2019learning, hafnerdream, hafner2021mastering, hafner2023mastering, hafner2025mastering]. While video generation tasks aim to maximize the visual fidelity of predicted sequences, robotic manipulation often does not require the full visual detail. Irrelevant elements such as the background or parts of the robot arm can be ignored. This motivates the use of object-centric representations, which focus on task-relevant entities and their interactions [ferraro2025focus, barcellona2025dream, villar2025playslot, zhi20253dflowaction, hong2024multiply]. Flow-centric representations, in contrast, are designed to capture the motion dynamics of the environment, emphasizing temporal change and spatial displacement [gao2024flip].
@@ III-D Task Scope
World models can also be categorized based on their task coverage. Some studies focus on single-task objectives, such as future-scene prediction [sudhakar2024controlling, finn2016unsupervised, barcellona2025dream, jang2025dreamgen, ebert2018robustness], or planning and action prediction [sekar2020planning].
In contrast, an increasing number of studies aim to support multiple tasks simultaneously, thereby enhancing the generality and applicability of world models. For instance, Cheang et al. [cheang2024gr, zhou2024robodreamer, du2023video, gao2024flip] generate videos for future-scene prediction and accordingly infer corresponding actions. Other works pursue simultaneous action prediction and world-scene forecasting [cen2025worldvla, chen2025egoagent, zhen20243d, song2025physical]. Beyond dual-task integration, several approaches extend world models to even broader capabilities. For instance, Bruce et al. [bruce2024genie] propose interactive video generation that supports environment prediction and imitation learning, and utilize a latent action model to infer policies from unseen, action-free videos. Liao et al. [liao2025genie] introduce a unified framework for embodied video generation, policy learning, and simulation. Lu et al. [lu2025gwm] learn 3D world representations for future-state prediction, imitation learning, and simulator through video generation. Zhu et al. [zhu2025irasim] develop an action-conditioned world model supporting trajectory-conditioned video generation, policy evaluation, and planning. Similarly, Huang et al. [huang2025enerverse] achieve multi-view video generation, robotic action prediction, and a data flywheel mechanism for sim-to-real adaptation.
Would Foundation Models. When discussing task scope, the notion of “foundation world models” becomes essential. These approaches aim to generalize across diverse tasks through large-scale training, paving the way for world models that act as universal backbones for robotics. One line of research achieves this through large-scale pretraining followed by task-specific fine-tuning [lu2025gwm, wu2024ivideogpt, agarwal2025cosmos, cheang2024gr, mazzaglia2024genrl]. In particular, Mazzaglia et al. [mazzaglia2024genrl] integrate a foundation VLM with a generative world model to enhance multimodal generalization. Other works directly pursue large-scale end-to-end training to build general-purpose world models [bruce2024genie, cen2025worldvla].
@@
IV Functions of World Models
World models play a central role in modern robotics by providing an internal predictive understanding of the environment. They enable robots to reason about future states, anticipate the consequences of actions, and perform evaluations, which are particularly valuable in real-world settings where interactions are costly, risky, or time-consuming. By modeling environmental dynamics, world models form the foundation for autonomous, adaptable, and efficient robotic systems. In robotic manipulation, world models serve two complementary functions: decision support, by predicting future scenes, actions and planning, and training facilitation, by generating data or acting as learned simulators. These roles are often closely related. For example, a world model used as a simulator can simultaneously generate training data and assist decision making [lu2025gwm, liao2025genie]. By combining these functionalities, world models provide a comprehensive framework that enables robots to act intelligently, learn efficiently, and adapt to complex and dynamic environments. Additional details are provided in Table I, which complements the following discussion.
@@ IV-A1 Implicit World Models for Action Prediction and Planning
This line of work explores world models that enable action prediction and planning without explicitly modeling state transitions or world dynamics. These approaches typically leverage the strong reasoning and next-token prediction capabilities of Large Language Models (LLMs), Vision-Language Models (VLMs), and Vision-Language-Action (VLA) models. Since LLMs lack direct access to environmental or robotic states, auxiliary components are often incorporated to provide grounding. For example, Ahn et al. [ahn2022can] introduce affordance functions to evaluate the feasibility of skills for completing a target task. Xiang et al. [xiang2024pandora, driess2023palm] employ encoders to process environmental information, while Zhang et al. [zhang2025dreamvla] integrate multimodal tokens including states, images, and text to enhance reasoning and generalization. Zhang et al. [huang2024embodied] further combine 2D and 3D encoders to process RGB images and 3D point clouds, capturing complementary spatial cues for richer world understanding. Hong et al. [hong2024multiply] extend this paradigm by incorporating additional sensory modalities such as vision, audio, tactile, and thermal inputs to achieve a more comprehensive understanding of the environment.
@@
Conventional LLMs are language-centric and typically treat visual and other sensory information as auxiliary inputs. VLMs extend this paradigm by jointly learning aligned visual and linguistic representations, enabling grounded perceptual understanding of the world [hu2023look]. Zhang et al. [zhao2024vlmpc] further leverage VLMs to generate candidate action sequences, which are evaluated using a lightweight action-conditioned video prediction model to forecast future scenes. The predicted outcomes are then assessed by the VLM to select the final action. An increasing number of studies extend VLMs to VLA by equipping them with low-level action generation capabilities. For instance, Zitkovich et al. [zitkovich2023rt] represent robot actions as a form of language, effectively bridging perception and control through textual grounding. Zhen et al. [zhen20243d] employ a 3D-based LLM [hong20233d] to represent and predict 3D world states and generate actions, incorporating a diffusion model to synthesize future scenes. Inspired by the dual-process theory of human cognition [kahneman2011thinking], Björck et al. [bjorck2025gr00t] design a dual-system architecture in which a VLM serves as the reasoning module (System 2) and a Diffusion Transformer functions as the action module (System 1), with both components jointly optimized for coordinated reasoning and actuation. Zhou et al. [zhou2025vision] preserve the reasoning capability of VLMs while introducing a Mixture-of-Experts to alleviate conflicts between multimodal understanding and robotic manipulation in the parameter space. Kim et al. [kim2025openvla] train their model on a large corpus of real-world robot demonstrations, enabling efficient adaptation to new robotic platforms through parameter-efficient fine-tuning.
@@
To further enhance long-horizon prediction, reasoning, and imagination, several methods integrate large language or multimodal models into other world model architectures, where they serve as core components. For instance, Chen et al. [chen2025egoagent] employ the open-source LLM, i.e., InternLM [team2023internlm], to predict future states from egocentric observations as a fundamental element of the JEPA framework. Similarly, Vision-Language Models [mazzaglia2024genrl] and Video-Language Models [wang2025founder] have been incorporated into Dreamer-style architectures for low-level dynamics modeling, where they extract high-level semantic knowledge of the world to guide prediction.
@@
Notably, LLMs, VLMs, and VLAs can also act as explicit world models that predict future scenes [zhang2025up, zhang2025dreamvla, zhao2025cot] or world knowledge [zhang2025dreamvla]. We will elaborate them in Section IV-A3.
@@ IV-A2 Latent Dynamics Modeling for Action Prediction and Planning
This line of research focuses on modeling the temporal evolution of environment dynamics within a latent space, facilitating efficient action prediction, planning and future imagination. Operating in a compact latent space requires fewer environment interactions and reduces computational cost compared to pixel-based modeling. Hafner et al. [hafner2019learning, hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering] introduce online planning in latent space through the Recurrent State-Space Model (RSSM), which learns to reconstruct input observations [hafner2019learning]. The Dreamer series [hafnerdream, hafner2021mastering, wu2023daydreamer, hafner2023mastering] introduces latent imagination, allowing agents to predict and plan over latent trajectories instead of pixels for more efficient and stable policy learning. Specifically, DreamerV1 [hafnerdream] learns long-horizon behaviors from images by jointly predicting actions and state values, greatly improving sample efficiency. DreamerV2 [hafner2021mastering] extends this framework to discrete environments by introducing binary latent variables, achieving human-level performance on the Atari benchmark. DreamerV3 [hafner2023mastering, hafner2025mastering] further improves scalability and generative capacity through techniques such as symlog normalization for reward stabilization, refined KL balancing, and enhanced replay buffers. Sekar et al. [sekar2020planning] enhance generalization to downstream tasks through self-supervised learning without task-specific rewards, while Wu et al. [wu2023daydreamer] deploy Dreamer in the real world without simulators. Gumbsch et al. [gumbsch2023learning] introduce context-sensitive dynamics via a context-specific RSSM and hierarchical architecture to improve scalability and long-horizon prediction. Ferraro et al. [ferraro2025focus] develop object-centric world models for improved interaction reasoning.
Under the JEPA framework, Chen et al. [chen2025egoagent] capture causal and temporal dependencies by organizing states and actions into an interleaved sequence, integrating future state prediction and action generation within a unified transformer architecture. Building on this, Assran et al. [assran2025v] leverage pre-trained video encoders optimized with a masked denoising objective as the core of JEPA, enabling self-supervised learning through an action-conditioned predictor that autoregressively forecasts future states and actions. Incorporating other potential world models, such as LLMs and VLMs, have been introduced in Section IV-A1.
There are also approaches that couple Model Predictive Control (MPC) with learned world models, where the predictive model is used to simulate future trajectories and select optimal actions in a receding-horizon manner. For example, Hansen et al. [hansen2022temporal] learn task-specific latent dynamics models using temporal-difference objectives and apply them for efficient online Model Predictive Control. Hansen et al. [hansen2024td] further improve generalization across diverse embodiments and action spaces by learning an implicit, control-centric dynamics model.
@@ IV-A3 Vision-based Action Prediction and Planning
Vision-Conditioned Goals. Finn et al. [finn2016unsupervised] learn to predict motion dynamics that remain consistent across visual appearances, aiming to enable long-range, action-conditioned video prediction and generalization to unseen objects. Ebert et al. [ebert2018robustness, ebert2018visual] improve long-horizon prediction by using an image registration–based cost function that continuously corrects errors during execution, achieving closed-loop visual planning. Bu et al. [bu2024closed] further extend this idea with text-conditioned video generation to synthesize depth- and flow-consistent sub-goal images. A feedback mechanism then selects sub-goals and generates corresponding actions based on visual error evaluation, bridging visual planning and policy learning.
@@
Imagining the future does not inherently produce actions. To enable action predictions, Finn et al. [finn2017deep, ebert2018robustness, ebert2018visual] incorporate visual prediction models with model-predictive control (MPC) to select the best action (sequence). Bu et al. [bu2024closed] use an error-measurement strategy to select the best sub-goal images and an MLP that is train with an Inverse Dynamics objective to decode the corresponding actions.
@@
Language-Conditioned Goals. In [finn2016unsupervised, finn2017deep, ebert2018robustness, ebert2018visual], task specifications are provided as goal images, which are often difficult to obtain and prone to over- or under-specification. To address this limitation, a growing line of research leverages language as a more flexible, compact, and general medium for specifying tasks. However, translating language instructions into precise, actionable representations grounded in the robot’s observations remains challenging due to the misalignment between linguistic descriptions and visual perception. To bridge this gap, Nair et al. [nair2022learning] use action-conditioned video prediction to simulate future scenes under different action sequences and learn a language-conditioned reward function from crowd-sourced descriptions to measure task completion. The best sequence is selected to maximize the reward. Zhang et al. [zhang2025up] take advantage of the semantic knowledge and reasoning abilities of VLA and incorporate a decoder into VLA to enable future scene predictions and action generation. Zhou et al. [zhou2024robodreamer] parse language instructions into compositional primitives to capture spatial object relations and generalize to novel commands, while also supporting multimodal task inputs such as goal images and sketches. Zhang et al. [zhang2025dreamvla] enhance reasoning and generalization by introducing additional dream queries that capture historical information and predict dynamic regions, depth, and semantic maps using foundation models such as DINOv2 [oquab2024dinov2] and SAM [kirillov2023segment].
@@
Diverse Goals. Some works leverage diverse goal conditions to improve task understanding and completion. For instance, Wang et al. [wang2025language] develop a language–gesture-conditioned video generation model to disambiguate task specifications and integrate a behavior-cloning policy that unifies visual plan generation and manipulation. Du et al. [du2023learning] incorporate observed images as additional context in each frame-denoising step to synthesize video plans and employ an inverse-dynamics model to infer the corresponding action sequences. Zhao et al. [zhao2025cot] introduce visual chain-of-thought reasoning into VLA models by autoregressively generating sub-goal images alongside language instructions, enabling temporal planning and improving reasoning capability.
@@
Action Inference. A key advantage of vision–based action prediction is that it does not rely on large-scale action-labeled data. They can be pre-trained on large-scale video data and infer actions by training a simple action extractor using small amounts of action data. Techniques for action extraction include the inverse dynamics model [liang2025video, du2023learning, zhou2024robodreamer], a transformer encoder-decoder architecture [wang2025language]. Zhang et al. [zhang2025combo] use vision language models to propose actions, and a tree search to find the best plan. However, video predictions would contain irrelevant information to the target tasks or actions to execute such as background and robot arm. To handle this, Zhi et al. [zhi20253dflowaction] extract 3D Flow from video data and learn 3D optical flow as a representation of object motions to guide action planning. Zhang et al. [zhang2025dreamvla] propose dynamic region-based forecasting, which leverages optical flow prediction model [karaev2024cotracker, karaev2024cotracker3] to identify dynamic regions within the scene, enabling the model to concentrate on areas of motion that are critical for task execution instead of redundant frame reconstruction. Agarwal et al. [agarwal2025cosmos] leverage large-scale pre-training on both image and post-training for robotic manipulation, including instruction-based video prediction and action-based next-frame prediction. 3D positional embeddings, including 3D factorized Rotary Position Embedding and absolute positional embedding for relative positions and absolute coordinates respectively, are adopted to capture spatial and temporal information. Actions are predicted through an action embedder MLP. Tian et al. [tian2025predictive] propose an end-to-end Predictive Inverse Dynamics Models, which learn actions and visual futures synergistically to enhance the simulation and action predictions ability. [guo2024prediction] predict both future frames and robot actions within joint latent denoising process, which support planning and acting in a closed-loop manner.
@@
Visual Fidelity vs. Action Prediction. Guo et al. [guo2025flowdreamer] hypothesize that models trained solely with frame-prediction losses tend to emphasize visual appearance fidelity while underestimating accurate dynamics modeling. This highlights the need for approaches that explicitly separate dynamics learning from visual rendering. To address this, FlowDreamer adopts a two-stage framework that first predicts environment dynamics and then renders corresponding visual observations.